A boxer is working out on the heavy bag. On one of the boxer's punches the 41 kg heavy bag goes from hanging at rest to swinging at 6.59 m/s. The boxer's glove is in contact with the heavy bag for only 0.15 seconds.
With how much force, in N, does the boxer hit the heavy bag?
1
Expert's answer
2020-12-14T12:13:45-0500

Impulse of the bag is:


Δp=mv0=mv\Delta p = mv-0 = mv

where m=41kgm = 41kg is the mass and v=6.59m/sv = 6.59 m/s is the final speed. here stands for the initial momentum (in rest).

The average force is (according to the Newton's second law):


N=ΔpΔtN = \dfrac{\Delta p}{\Delta t}

where Δt=0.15s\Delta t = 0.15s is the time of interaction. Thus, obtain:


F=mvΔtF=41kg6.59m/s0.15s1801NF = \dfrac{mv}{\Delta t}\\ F= \dfrac{ 41kg\cdot 6.59m/s }{0.15s} \approx 1801N

Answer. 1801 N.